# Clairaut anti-invariant Riemannian map from/to Kähler manifolds admitting Ricci solitons

### v1

@@ 4 Clairaut Anti-invariant Riemannian maps to Kähler manifolds admitting Ricci Solitons
$\begin{split}Ric(QD,QE)=Ric^{(rangeF_{*})^{\perp}}(QD,QE)-(m-r)\Big{(}QD(g)QE(g)+g_{N}(\nabla^{M}_{QD}\nabla^{M}g,QE)\Big{)}.\end{split}$ (4.9)
@@
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut Langragian Riemannian map from a Riemannian manifold to a Kähler manifold with $\tilde{s}=e^{g}$. Then Ricci curvature is given by
$Ric(F_{*}X,F_{*}Y)=Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{JF_{*}X}JF_{*}Y)$ (4.10)
$\begin{split}Ric(F_{*}X,E)=&\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{*_{F_{*}PE}}S)_{JF_{*}X}F_{*}X_{j},F_{*}X_{j})\\ &-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{JF_{*}X}PE,F_{*}X_{j})\\ &-\sum_{k=1}^{r_{1}}g_{M}(R^{F\perp}(PE,e_{k})JF_{*}X,e_{k}),\end{split}$ (4.11)
$\begin{split}Ric(D,E)=Ric^{(rangeF_{*})}(PD,PE)-\sum_{k=1}^{r_{1}}g_{M}(PD,PE)\Big{(}(||\nabla^{M}g||^{2}+H^{g}(e_{k},e_{k})\Big{)},\end{split}$ (4.12)
@@
In case of Langrangian Riemannian map, $\nu=0,$ this implies $QD=QE=0.$ Applying this in (4.1), (4.2) and (4.3), we get the required result. ∎
@@
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with vertical potential vector field $\vartheta$ and scalar $\lambda$ such that $(rangeF_{*})^{\perp}$ is an Einstein. Then $\vartheta$ is a conformal vector field on $rangeF_{*}$.
@@
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with vertical potential vector field $\vartheta\in\Gamma(TM)$, we have
$\frac{1}{2}(L_{\vartheta}g_{M})(F_{*}X,F_{*}Y)+Ric(F_{*}X,F_{*}Y)+\lambda g_{M}(F_{*}X,F_{*}Y)=0.$
@@
$\begin{split}&\frac{1}{2}(L_{\vartheta}g_{M})(F_{*}X,F_{*}Y)+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{JF_{*}X}JF_{*}Y)+\lambda g_{M}(F_{*}X,F_{*}Y)=0,\end{split}$
@@
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla_{F_{*}X}\vartheta,F_{*}Y)+g_{M}(\nabla_{F_{*}Y}\vartheta,F_{*}X)\Big{)}+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{JF_{*}X}JF_{*}Y)+\lambda g_{M}(F_{*}X,F_{*}Y)=0.\end{split}$
Since $F$ is a totally geodesic Clairaut anti-invariant Riemannian map, second fundamental form of $F$ vanishes. Then by Theorem 3.2 ([7]), g is constant on $M.$ Therefore, we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla^{N}_{F_{*}X}\vartheta,F_{*}Y)+g_{M}(\nabla^{N}_{F_{*}Y}\vartheta,F_{*}X)\Big{)}+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+\lambda g_{M}(F_{*}X,F_{*}Y)=0.\end{split}$ (4.13)
@@
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla^{M}_{F_{*}X}F_{*}Z,F_{*}Y)+g_{M}(\nabla^{M}_{F_{*}Y}F_{*}Z,F_{*}X)\Big{)}+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+\lambda g_{M}(JF_{*}X,JF_{*}Y)=0.\end{split}$ (4.14)
@@
$Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)=\lambda g_{M}(JF_{*}X,JF_{*}Y)=\lambda g_{M}(F_{*}X,F_{*}Y),$
@@
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla^{M}_{F_{*}X}F_{*}Z,F_{*}Y)+g_{M}(\nabla^{M}_{F_{*}Y}F_{*}Z,F_{*}X)\Big{)}+2\lambda g_{M}(F_{*}X,F_{*}Y)=0,\end{split}$
@@
$\frac{1}{2}(L_{F_{*}Z}g_{M})(F_{*}X,F_{*}Y)+2\lambda g_{M}(F_{*}X,F_{*}Y)=0.$
This implies that $F_{*}Z$ is conformal vector field on $(rangeF_{*}).$
@@
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with horizontal potential vector field $\vartheta$ and scalar $\lambda.$ Then $J(rangeF_{*})$ is an Einstein.
@@
Since $\vartheta$ is horizontal vector field, i.e., $\vartheta=V,$ then from (4.13), we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla^{M}_{F_{*}X}V,F_{*}Y)+g_{M}(\nabla^{M}_{F_{*}Y}V,F_{*}X)\Big{)}+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+\lambda g_{M}(F_{*}X,F_{*}Y)=0.\end{split}$
Using (2.4) and Theorem 3.1([7]), we get
$V(g)g_{M}(F_{*}X,F_{*}Y)+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)+\lambda g_{M}(JF_{*}X,JF_{*}Y)=0.$
This implies
$Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)+\rho g_{M}(JF_{*}X,JF_{*}Y)=0,$
where $\rho=V(g)+\lambda.$ This proves that $J(rangeF_{*})$ is an Einstein. ∎
@@
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut Langragian Riemannian map from a Riemannian manfold $(N,g_{N})$ to a Kähler manifold $(M,g_{M},J),$ admitting Ricci soliton with horizontal potential vector field $\vartheta$ and scalar $\lambda$ then, $(rangeF_{*})^{\perp}$ is an Einstein.
@@
Since $F$ is a Langragian Riemannian map, $J(rangeF_{*})=(rangeF_{*})^{\perp}.$ Using Theorem (4.3), we see that $(rangeF_{*})^{\perp}$ is an Einstein. ∎
@@
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut Langragian Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with horizontal potential vector field $\vartheta\in\Gamma(TM)$ and scalar $\lambda.$ Then scalar curvature $s^{(rangeF_{*})^{\perp}}$ of $(rangeF_{*})^{\perp}$ is given by
$s^{(rangeF_{*})^{\perp}}=-\rho(m-r),$ (4.15)
where $\rho=V(g)+\lambda.$
@@
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with horizontal potential vector field $\vartheta\in\Gamma(TM)$, we have
$\frac{1}{2}(L_{\vartheta}g_{M})(F_{*}X,F_{*}Y)+Ric(F_{*}X,F_{*}Y)+\lambda g_{M}(F_{*}X,F_{*}Y)=0.$
@@
$\begin{split}&\frac{1}{2}(L_{\vartheta}g_{M})(F_{*}X,F_{*}Y)+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{JF_{*}X}JF_{*}Y)+\lambda g_{M}(F_{*}X,F_{*}Y)=0.\end{split}$
In Theorem 4.3, we have seen that $\frac{1}{2}(L_{\vartheta}g_{M})(F_{*}X,F_{*}Y)$ = $V(g)g_{M}(F_{*}X,F_{*}Y),$ using Theorem 5.2 [7], we get
$\begin{split}&Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)+\rho g_{M}(F_{*}X,F_{*}Y)=0,\end{split}$
@@
$\begin{split}Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)=-\rho g_{M}(JF_{*}X,JF_{*}Y).\end{split}$
Taking trace of above equation, we get
$\begin{split}s^{(rangeF_{*})^{\perp}}=-\rho n_{1}\end{split}$
This completes the proof. ∎
Theorem 4.5.
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut Langangian Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with vertical potential vector field $\vartheta$ and scalar $\lambda.$ Then $rangeF_{*}$ is an Einstein.
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with vertical potential vector field $\vartheta\in\Gamma(TM)$, we have
$\frac{1}{2}(L_{\vartheta}g_{M})(D,E)+Ric(D,E)+\lambda g_{M}(D,E)=0.$
Using (4.12) in above equation, we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla_{D}\vartheta,E)+g_{M}(\nabla_{E}\vartheta,D)+Ric^{(rangeF_{*})}(PD,PE)\\ &+\sum_{k=1}^{r_{1}}g_{M}(PD,PE)\Big{(}-H^{g}(e_{k},e_{k})-||\nabla^{M}g||^{2}+\lambda\Big{)}=0.\end{split}$ (4.16)
Since $\vartheta$ is vertical, i.e., $\vartheta=F_{*}X,$ we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla_{D}F_{*}X,E)+g_{M}(\nabla_{D}F_{*}X,E)+Ric^{(rangeF_{*})}(PD,PE)\\ &+\sum_{k=1}^{r_{1}}g_{M}(PD,PE)(\lambda-H^{g}(e_{k},e_{k})-||\nabla^{M}g||^{2})=0.\end{split}$
Using metric compatibility, we obtain
$Ric^{(rangeF_{*})}+\psi g_{M}(PD,PE)=0,$ (4.17)
where $\psi=\sum_{k=1}^{r_{1}}(\lambda-H^{g}(e_{k},e_{k})-||\nabla^{M}g||^{2})$ is a function on $M$. This shows that $rangeF_{*}$ is an Einstein. ∎
Theorem 4.6.
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut Langangian Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with horizontal potential vector field $\vartheta$ and a scalar $\lambda.$ Then the scalar curvature $s^{(rangeF_{*})^{\perp}}$ of $(rangeF_{*})^{\perp}$ is given by $s^{(rangeF_{*})^{\perp}}=-r\psi.$
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with horizontal potential vector field $\vartheta\in\Gamma(TM)$, we have
$\begin{split}\frac{1}{2}(L_{\vartheta}g_{M})(D,E)+Ric(D,E)+\lambda g_{M}(D,E)=0.\end{split}$
Using (4.12) and since $\vartheta$ is horizontal vector field, i.e., $\vartheta=W,$ we have
$\begin{split}\frac{1}{2}\Big{(}g_{M}(\nabla_{D}W,E)+g_{M}(D,\nabla_{E}W)\Big{)}+Ric^{(rangeF_{*})}(PD,PE)+\psi g_{M}(PD,PE)=0.\end{split}$
Taking trace of above equation, we get
$\sum_{k=1}^{r_{1}}g_{M}(\nabla_{e_{k}}e_{k},e_{k})+\sum_{j=r+1}^{m}Ric^{(rangeF_{*})}(F_{*}{X_{j}},F_{*}{X_{j}})+\sum_{j=r+1}^{m}\psi g_{M}(F_{*}{X_{j}},F_{*}{X_{j}})=0$
which gives
$s^{(rangeF_{*})}=-r\psi.$ (4.18)
@@
Let $M=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}):x_{i}\neq 0,i=1,...,6\}$ be a Riemannian manifold with Riemannian metric $g_{M}=dx_{1}^{2}+dx_{2}^{2}+dx_{3}^{2}+dx_{4}^{2}+dx_{5}^{2}+dx_{6}^{2}$ and $N=R^{6}$ be a Kähler manifold with Riemannian metric and complex structure $g_{N}=e^{-2y_{4}}dy_{1}^{2}+dy_{2}^{2}+e^{-2y_{4}}dy_{3}^{2}+dy_{4}^{2}+dy_{5}^{2}+dy_{6}^{2},$ $J(a,b,c,d,l,h)=(-b,a,-d,c,-h,l)$ respectively. We construct a map $F:(M,g_{M})\rightarrow(N,g_{N},J),$ defined by
$F(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})=\big{(}\frac{x_{1}-x_{4}}{\sqrt{2}},0,x_{2},0,0,0\Big{)}$
Then $kerF_{*}=span\{U_{1}=\frac{e_{1}+e_{4}}{\sqrt{2}},U_{2}=e_{3},U_{3}=e_{5},U_{4}=e_{6}\}$ and $(kerF_{*})^{\perp}=span\{X_{1}=\frac{e_{1}-e_{4}}{\sqrt{2}},X_{2}=e_{2}\},$ where $\{e_{1}=\frac{\partial}{\partial x_{1}},e_{2}=\frac{\partial}{\partial x_{2}},e_{3}=\frac{\partial}{\partial x_{3}},e_{4}=\frac{\partial}{\partial x_{4}},e_{5}=\frac{\partial}{\partial x_{5}},e_{6}=\frac{\partial}{\partial x_{6}}\}$ and $\{e_{1}^{\prime}=e^{y_{4}}\frac{\partial}{\partial y_{1}},e_{2}^{\prime}=\frac{\partial}{\partial y_{2}},e_{3}^{\prime}=e^{y_{4}}\frac{\partial}{\partial y_{3}},e_{4}^{\prime}=\frac{\partial}{\partial y_{4}},e_{5}^{\prime}=\frac{\partial}{\partial y_{5}},e_{6}^{\prime}=\frac{\partial}{\partial y_{6}}\}$ are the bases of $T_{p}M$ and $T_{F(p)N}$ respectively.
Here $F_{*}(X_{1})=e_{1}^{\prime}$ and $F_{*}(X_{2})=e_{3}^{\prime}$ and we can easily compute that

### v2

@@ 4 Clairaut Anti-invariant Riemannian maps to Kähler manifolds admitting Ricci Solitons
$\begin{split}Ric(QD,QE)=Ric^{(rangeF_{*})^{\perp}}(QD,QE)-(m-r)\Big{(}QD(g)QE(g)+g_{N}(\nabla^{N}_{QD}\nabla^{N}g,QE)\Big{)}.\end{split}$ (4.9)
@@
Let $F:(M,g_{M})\rightarrow(N,g_{N},J^{\prime})$ be a Clairaut Lagrangian Riemannian map from a Riemannian manifold to a Kähler manifold with $\tilde{s}=e^{g}$. Then Ricci curvature is given by
$Ric(F_{*}X,F_{*}Y)=Ric^{(rangeF_{*})^{\perp}}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)$ (4.10)
$\begin{split}Ric(F_{*}X,E)=&\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{{}^{*}_{F_{*}PE}}S)_{J^{\prime}F_{*}X}F_{*}X_{j},F_{*}X_{j})\\ &-\sum_{j=r+1}^{m}g_{N}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{J^{\prime}F_{*}X}PE,F_{*}X_{j})\\ &-\sum_{k=1}^{r_{1}}g_{N}(R^{F\perp}(PE,e_{k})J^{\prime}F_{*}X,e_{k}),\end{split}$ (4.11)
$\begin{split}Ric(D,E)=Ric^{(rangeF_{*})}(PD,PE)\end{split}$ (4.12)
@@
In case of Lagrangian Riemannian map, $\nu=0,$ this implies $QD=QE=0.$ From Theorem 5.2 of [11], we have $f$ is constant on $J^{\prime}(rangeF_{*}).$ Applying these things in (4.1), (4.2) and (4.3), we get the required result. ∎
@@
Let $F:(M,g_{M})\rightarrow(N,g_{N},J^{\prime})$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with vertical potential vector field $\vartheta$ and scalar $\lambda$ such that $(rangeF_{*})^{\perp}$ is an Einstein. Then $\vartheta$ is a conformal vector field on $rangeF_{*}$.
@@
Since $(N,g_{N},J^{\prime})$ is a Kähler manifold admitting a Ricci soliton with vertical potential vector field $\vartheta\in\Gamma(TN)$, we have
$\frac{1}{2}(L_{\vartheta}g_{N})(F_{*}X,F_{*}Y)+Ric(F_{*}X,F_{*}Y)+\lambda g_{N}(F_{*}X,F_{*}Y)=0.$
@@
$\begin{split}&\frac{1}{2}(L_{\vartheta}g_{N})(F_{*}X,F_{*}Y)+Ric^{(rangeF_{*})^{\perp}}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)\\ &+(m-r)g_{N}(\nabla^{N}g,\nabla^{F\perp}_{J^{\prime}F_{*}X}J^{\prime}F_{*}Y)+\lambda g_{N}(F_{*}X,F_{*}Y)=0,\end{split}$
@@
$\begin{split}&\frac{1}{2}\Big{(}g_{N}(\nabla_{F_{*}X}\vartheta,F_{*}Y)+g_{N}(\nabla_{F_{*}Y}\vartheta,F_{*}X)\Big{)}+Ric^{(rangeF_{*})^{\perp}}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)\\ &+(m-r)g_{N}(\nabla^{N}g,\nabla^{F\perp}_{J^{\prime}F_{*}X}J^{\prime}F_{*}Y)+\lambda g_{N}(F_{*}X,F_{*}Y)=0.\end{split}$
Since $F$ is a totally geodesic Clairaut anti-invariant Riemannian map, second fundamental form of $F$ vanishes. Then by Theorem 3.2 of [11], g is constant on $N.$ Therefore, we get
$\begin{split}&\frac{1}{2}\Big{(}g_{N}(\nabla^{N}_{F_{*}X}\vartheta,F_{*}Y)+g_{N}(\nabla^{N}_{F_{*}Y}\vartheta,F_{*}X)\Big{)}+Ric^{(rangeF_{*})^{\perp}}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)\\ &+\lambda g_{N}(F_{*}X,F_{*}Y)=0.\end{split}$ (4.13)
@@
$\begin{split}&\frac{1}{2}\Big{(}g_{N}(\nabla^{N}_{F_{*}X}F_{*}Z,F_{*}Y)+g_{N}(\nabla^{N}_{F_{*}Y}F_{*}Z,F_{*}X)\Big{)}+Ric^{(rangeF_{*})^{\perp}}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)\\ &+\lambda g_{N}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)=0.\end{split}$ (4.14)
@@
$Ric^{(rangeF_{*})^{\perp}}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)=\lambda g_{N}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)=\lambda g_{N}(F_{*}X,F_{*}Y),$
@@
$\begin{split}&\frac{1}{2}\Big{(}g_{N}(\nabla^{N}_{F_{*}X}F_{*}Z,F_{*}Y)+g_{N}(\nabla^{N}_{F_{*}Y}F_{*}Z,F_{*}X)\Big{)}+2\lambda g_{N}(F_{*}X,F_{*}Y)=0,\end{split}$
@@
$\frac{1}{2}(L_{F_{*}Z}g_{N})(F_{*}X,F_{*}Y)+2\lambda g_{N}(F_{*}X,F_{*}Y)=0.$
This implies that $F_{*}Z$ is a conformal vector field on $(rangeF_{*}).$
@@
Let $F:(M,g_{M})\rightarrow(N,g_{N},J^{\prime})$ be a Clairaut Lagrangian Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with horizontal potential vector field $\vartheta$ and scalar $\lambda.$ Then $(rangeF_{*})^{\perp}$ is an Einstein.
@@
Since $(N,g_{N},J^{\prime})$ is a Kähler manifold admitting a Ricci soliton with horizontal potential vector field $\vartheta\in\Gamma(TN)$, we have
$\frac{1}{2}(L_{\vartheta}g_{N})(F_{*}X,F_{*}Y)+Ric(F_{*}X,F_{*}Y)+\lambda g_{N}(F_{*}X,F_{*}Y)=0.$
Using (4.10) in above equation, we get
$\begin{split}&\frac{1}{2}(L_{\vartheta}g_{N})(F_{*}X,F_{*}Y)+Ric^{(rangeF_{*})^{\perp}}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)\\ &+\lambda g_{N}(F_{*}X,F_{*}Y)=0.\end{split}$
Since $\vartheta$ is horizontal, i.e., $\vartheta=D,$ with the help of (2.4), and Theorem 3.1 of [11], we get
$D(g)g_{N}(F_{*}X,F_{*}Y)+Ric^{(rangeF_{*})^{\perp}}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)+\lambda g_{N}(F_{*}X,F_{*}Y)=0.$ (4.15)
Since $N$ is a Kähler manifold, we have
$g_{N}(F_{*}X,F_{*}Y)=g_{N}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y),$
then applying above equation in (4.15), we get
$Ric^{(rangeF_{*})^{\perp}}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)+\rho g_{M}(J^{\prime}F_{*}X,J^{\prime}F_{*}Y)=0,$ (4.16)
where $\rho=D(g)+\lambda.$ This proves that $(rangeF_{*})^{\perp}$ is an Einstein. ∎
@@
Let $F:(M,g_{M})\rightarrow(N,g_{N},J^{\prime})$ be a Clairaut Lagrangian Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with horizontal potential vector field $\vartheta\in\Gamma(TN)$ and scalar $\lambda.$ Then scalar curvature $s^{(rangeF_{*})^{\perp}}$ of $(rangeF_{*})^{\perp}$ is given by
$s^{(rangeF_{*})^{\perp}}=-\rho n_{1},$ (4.17)
where $\rho=D(g)+\lambda.$
@@
Taking trace of (4.16), we get the required result. ∎
@@
Let $F:(M,g_{M})\rightarrow(N,g_{N},J^{\prime})$ be a Clairaut Lagrangian Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with vertical potential vector field $\vartheta$ and scalar $\lambda.$ Then $rangeF_{*}$ is an Einstein.
@@
Since $(N,g_{N},J^{\prime})$ is a Kähler manifold admitting a Ricci soliton with vertical potential vector field $\vartheta\in\Gamma(TN)$, we have
$\frac{1}{2}(L_{\vartheta}g_{N})(D,E)+Ric(D,E)+\lambda g_{N}(D,E)=0.$
@@
$\begin{split}&\frac{1}{2}\Big{(}g_{N}(\nabla_{D}\vartheta,E)+g_{N}(\nabla_{E}\vartheta,D)+Ric^{(rangeF_{*})}(PD,PE)\\ &+\lambda g_{N}(PD,PE)=0.\end{split}$ (4.18)
Since $\vartheta$ is vertical, i.e., $\vartheta=F_{*}X,$ we get
$\begin{split}&\frac{1}{2}\Big{(}g_{N}(\nabla_{D}F_{*}X,E)+g_{N}(\nabla_{E}F_{*}X,D)+Ric^{(rangeF_{*})}(PD,PE)\\ &+\lambda g_{N}(PD,PE)=0.\end{split}$
Using metric compatibility, we obtain
$Ric^{(rangeF_{*})}(PD,PE)+\lambda g_{N}(PD,PE)=0,$ (4.19)
This shows that $rangeF_{*}$ is an Einstein. ∎
Theorem 4.5.
Let $F:(M,g_{M})\rightarrow(N,g_{N},J^{\prime})$ be a Clairaut Lagrangian Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with horizontal potential vector field $\vartheta$ and a scalar $\lambda.$ Then the scalar curvature $s^{(rangeF_{*})}$ of $(rangeF_{*})$ is given by $s^{(rangeF_{*})}=-(m-r)\lambda.$
Proof.
Since $(N,g_{N},J^{\prime})$ is a Kähler manifold admitting a Ricci soliton with horizontal potential vector field $\vartheta\in\Gamma(TN)$, we have
$\begin{split}\frac{1}{2}(L_{\vartheta}g_{N})(D,E)+Ric(D,E)+\lambda g_{N}(D,E)=0.\end{split}$
Since $\vartheta$ is horizontal vector field and using (4.12), we have
$\begin{split}\frac{1}{2}\Big{(}g_{N}(\nabla_{D}\vartheta,E)+g_{N}(D,\nabla_{E}\vartheta)\Big{)}+Ric^{(rangeF_{*})}(PD,PE)+\lambda g_{N}(PD,PE)=0.\end{split}$
Taking trace of above equation, we get
$\sum_{k=1}^{r_{1}}g_{N}(\nabla_{e_{k}}e_{k},e_{k})+\sum_{j=r+1}^{m}Ric^{(rangeF_{*})}(F_{*}{X_{j}},F_{*}{X_{j}})+\sum_{j=r+1}^{m}\lambda g_{N}(F_{*}{X_{j}},F_{*}{X_{j}})=0$
@@
$s^{(rangeF_{*})}=-(m-r)\lambda.$ (4.20)
@@
Let $M=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}):x_{i}\neq 0,i=1,...,6\}$ be a Riemannian manifold with Riemannian metric $g_{M}=dx_{1}^{2}+dx_{2}^{2}+e^{2x_{5}}dx_{3}^{2}+e^{2x_{5}}dx_{4}^{2}+dx_{5}^{2}+e^{2x_{5}}dx_{6}^{2}$ and $N=R^{6}$ be a Kähler manifold with Riemannian metric and complex structure $g_{N}=dy_{1}^{2}+dy_{2}^{2}+e^{2y_{5}}dy_{3}^{2}+dy_{4}^{2}+dy_{5}^{2}+e^{2y_{5}}dy_{6}^{2},$ $J^{\prime}(a,b,c,d,l,h)=(-b,a,-d,c,-h,l)$ respectively. We construct a map $F:(M,g_{M})\rightarrow(N,g_{N},J^{\prime}),$ defined by
$F(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})=\big{(}0,x_{2},0,0,x_{5},0\Big{)}$
Then $kerF_{*}=span\{U_{1}=e_{1},U_{2}=e_{3},U_{3}=e_{4},U_{4}=e_{6}\}$ and $(kerF_{*})^{\perp}=span\{X_{1}=e_{2},X_{2}=e_{5}\},$ where $\{e_{1}=\frac{\partial}{\partial x_{1}},e_{2}=\frac{\partial}{\partial x_{2}},e_{3}=e^{-x_{5}}\frac{\partial}{\partial x_{3}},e_{4}=e^{-x_{5}}\frac{\partial}{\partial x_{4}},e_{5}=\frac{\partial}{\partial x_{5}},e_{6}=e^{-x_{5}}\frac{\partial}{\partial x_{6}}\}$ and $\{e_{1}^{\prime}=\frac{\partial}{\partial y_{1}},e_{2}^{\prime}=\frac{\partial}{\partial y_{2}},e_{3}^{\prime}=e^{-y_{5}}\frac{\partial}{\partial y_{3}},e_{4}^{\prime}=\frac{\partial}{\partial y_{4}},e_{5}^{\prime}=\frac{\partial}{\partial y_{5}},e_{6}^{\prime}=e^{-y_{5}}\frac{\partial}{\partial y_{6}}\}$ are the bases of $T_{p}M$ and $T_{F(p)N}$ respectively.
Here $F_{*}(X_{1})=e_{2}^{\prime}$ and $F_{*}(X_{2})=e_{5}^{\prime}$ and we can easily compute that